OpenStudy (anonymous):

Factor the expression: x^2 + 3x + 2 does this mean just combine like factors?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x^2 + 3x + 2 --> We need to find numbers that multiply to 2 and add up to positive 3

OpenStudy (anonymous):

oml lol who invented math -.-

OpenStudy (anonymous):

ahem.. 2 * 1 lol

OpenStudy (anonymous):

bahaha wait so the answer is just 2?

OpenStudy (anonymous):

so the factors are: (x + 2)(x + 1)

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

you seriously have never factored? lol

OpenStudy (anonymous):

that is a slight issue lol are you sure you understand that steps?

OpenStudy (anonymous):

Nm dont read that lol

OpenStudy (anonymous):

so it would be x^2+3x+2?

OpenStudy (anonymous):

no we are factoring x^2 + 3x + 2 --> the factors are (x+2)(x+1)

OpenStudy (anonymous):

Ohhh ok i understand!

OpenStudy (mathstudent55):

Factoring means undoing a multiplication. For example, take the number 15. Factoring 15 into its prime factors means show which prime numbers have a product of 15. 15 = 3 * 5, so the prime factorization of 15 is 3 * 5. Now let's look at at an expression with a variable, x. Factor \(x^2 + x\) You need to find two factors that multiply to \(x^2 + x\) Since the terms x^2 and x have a common factor, x, you can factor that common factor out: \(x^2 + x = x(x + 1)\) This is the factorization of \(x^2 + x\). Then you can get into factoring expressions like yours. You have a trinomial, a three-term polynomial, with a squared variable. \(x^2 + 3x + 2\) In this case, you look for two number that multiply to the last number, 2, and add to the number of the middle term, 3. In this case the numbers are 2 and 1, since 2 * 1 = 2, and 2 + 1 = 3. You write the factorization as \(x^2 + 3x + 2 = (x = 2)(x + 1)\)